OpenStudy (anonymous):

Kim bought 4 shirts and 3 pairs of jeans for $109.85. Jim bought 6 shirts and 1 pair of jeans for $94.95. Each shirt costs the same amount. Each pair of jeans costs the same amount. What is the cost, in dollars, for 1 pair of jeans?

OpenStudy (mathstudent55):

Write a system of equations, then solve it. Let number of shirts = s Let number of jeans = j Kim: 4s + 3j = 109.85 Jim: 6s + j = 94.95 The system of equations is: 4s + 3j = 109.85 6s + j = 94.95 Do you know how to solve a system of equations?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

ok, great

OpenStudy (anonymous):

so am I solving for j then?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes.....

OpenStudy (anonymous):

but how does that give me the cost of the jeans?

OpenStudy (anonymous):

the j is representing the cost of 1 pair of jeans not the number.......the number of them are the numbers given that are multiplying the j.......

OpenStudy (mathstudent55):

Solve the second equation for j, which is easy. Then substitute that into the first equation. Then you solve the first equation for s. Once you have s, insert it into either of the two original equations and solve for j.

OpenStudy (anonymous):

j=-6s+94.95

OpenStudy (mathstudent55):

4s + 3j = 109.85 6s + j = 94.95 ----> j = 94.95 - 6s Now that you have the second equation solved for j, insert what j is into j of the first equation. Yes you got j correct.

OpenStudy (anonymous):

4s+3(-6s+94.95)=109.85

OpenStudy (mathstudent55):

First equation: 4s + 3j = 109.85 now substitute j in this eq with -6s + 94.95

OpenStudy (mathstudent55):

excellent, you're doing it before I even write it

OpenStudy (mathstudent55):

Now multiply out the 3 on the left side

OpenStudy (anonymous):

s=12.71 ?

OpenStudy (mathstudent55):

4s+3(-6s+94.95)=109.85 4s + 3*(-6s) + 3*94.95) = 109.85 4s - 18s + 284.85 = 109.85 -14s = -175 s = 12.5

OpenStudy (anonymous):

s= 12.50

OpenStudy (anonymous):

Ya, I put -178 instead of -175

OpenStudy (anonymous):

THANKS!!!!

OpenStudy (mathstudent55):

Now replace s = 12.5 into either of the original equations to get j.

OpenStudy (mathstudent55):

Are you sure you copied the problem correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is $12.50 for the shirts

OpenStudy (anonymous):

j would equal $19.95

OpenStudy (mathstudent55):

Right. $12.50 is the price, not the number of shirts. It does make sense.

OpenStudy (mathstudent55):

Correct. Good job @ballerina789

OpenStudy (anonymous):

I just needed a boost. Thanks for that.